# U-index: A Universal Indexing Framework for Matching Long Patterns

### v1

@@ Abstract
Motivation. Text indexing is a fundamental and well-studied problem. Classic solutions to this problem either replace the original text with a compressed representation, e.g., the FM-index and its variants, or keep it uncompressed but attach some redundancy — an index — to accelerate matching, e.g., the suffix array. The former solutions thus retain excellent compressed space, but are practically slow to construct and query. The latter approaches, instead, sacrifice space efficiency but are typically faster; for example, the suffix array takes much more space than the text itself for commonly used alphabets, like ASCII or DNA, but it is fast to construct and query.
Methods. In this paper, we show that efficient text indexing can be achieved using just a small extra space on top of the original text, provided that the query patterns are sufficiently long. More specifically, we develop a new indexing paradigm in which a sketch of a query pattern is first matched against a sketch of the text. Once candidate matches are retrieved, they are verified using the original text. This paradigm is thus universal in the sense that it allows us to use any solution to index the sketched text, like a suffix array, FM-index, or r-index.
Results. We explore both the theory and the practice of this universal framework. With an extensive experimental analysis, we show that, surprisingly, universal indexes can be constructed much faster than their unsketched counterparts and take a fraction of the space, as a direct consequence of (i) having a lower bound on the length of patterns and (ii) working in sketch space. Furthermore, these data structures have the potential of retaining or even improving query time, because matching against the sketched text is faster and verifying candidates can be theoretically done in constant time per occurrence (or, in practice, by short and cache-friendly scans of the text).
Finally, we discuss some important applications of this novel indexing paradigm to computational biology. We hypothesize that such indexes will be particularly effective when the queries are sufficiently long, and so demonstrate applications in long-read mapping.
@@
Sketching; Hashing; Minimizers; Text Indexing
@@ 1 Introduction
Given a string $T[0\mathinner{.\,.}n)$ (the “text” henceforth) of $n$ characters over an alphabet $\Sigma=[0,\sigma)$, we are asked to pre-process $T$ so that the following queries are supported efficiently for any pattern $P[0\mathinner{.\,.}m)$ and any $0\leq i<j\leq n$:
@@
On the other hand, uncompressed approaches, such as the suffix array [28], do not alter the original text but instead attach some form of redundancy — henceforth referred to as the index — to accelerate pattern matching. These methods, while being faster in terms of query times, often suffer from significant space inefficiency. For example, suffix arrays usually take up more space than the text itself, particularly for commonly used alphabets such as ASCII and DNA sequences [20]. This makes such structures impractical for many applications, especially when working with very large datasets.
Our Contributions. In this paper, we focus on the latter scenario of uncompressed text indexing and aim to address its space inefficiency while retaining efficient queries. In the following, we discuss the Locate query only, given that Count can be trivially implemented using Locate, and Extract is done by explicitly accessing the text $T$. We propose a novel approach that solves the text indexing problem using only a small amount of additional space beyond the original text, without sacrificing the speed of query processing. This is achieved under the assumption that the query patterns $P[0\mathinner{.\,.}m)$ are sufficiently long, i.e., we require that $m\geq\ell$ for some fixed lower bound $\ell>0$ [4], and operate under the assumption that $n\gg\ell$. A typical value for $\ell$ lies in $[32,1000]$. The core idea of our approach is to transform the text into sketch space, and to construct and use a smaller index over the sketched text to enable queries with only small additional space. Generally speaking, a sketch is a compact representation of an object that retains enough information for approximate matches.
@@
    Each query pattern $P$ is then sketched into $Q=\textup{{Sketch}}(P)$, and $Q$ is matched against $\textup{{Index}}(S)$ to identify potential matches $L^{\prime}=\textup{{Locate}}(S,Q)$.
@@
    Candidate matches in $L^{\prime}$ are mapped back to their positions in $T$, where we verify that indeed the query pattern $P$ matches the text. Thus we obtain $L=\textup{{Map}}(L^{\prime})$, where $L=\textup{{Locate}}(P,T)$.
The universality of our framework lies in its flexibility: (i) any indexing structure Index, such as the FM-index or the suffix array, can be used to index the sketched text $S$, making the method adaptable to a variety of text indexing techniques [28, 13, 19, 14]; and (ii) any locally consistent sampling mechanism, Sketch, such as minimizers [31, 32], syncmers [10], or bd-anchors [27], can be used to sketch $S$ and $P$ making the method adaptable to a variety of sampling techniques. We also expect that our framework, which relies on $\textup{{Index}}(S)$, has competitive queries compared to the ones of $\textup{{Index}}(T)$ or possibly faster when the query time of $\textup{{Index}}(T)$ is a function of $n=|T|$. This is merely because $|S|<|T|$ (in practice, depending on the sketching mechanism used, $|S|$ could be $\Theta(|T|/\ell)$ or $\Theta(|T|/(\ell-k))$ for some $k=o(\ell)$). A typical such case is when Index is the suffix array [28] and binary search will take place on a smaller array, i.e., the suffix array of $S$.
We explore both the theoretical underpinnings and the practical implementation of this universal framework. We demonstrate that universal indexes can be constructed significantly faster and occupy a fraction of the space compared to their unsketched counterparts. These space savings are a direct consequence of operating in the reduced “sketch space”. Depending on the index used, the performance of pattern matching in sketched space can be either faster (suffix array) or slower (FM-index). In either case the overall query performance is slightly slower though, due to a relatively large number of false positive matches that have to be rejected. This verification step can be performed in constant time per occurrence (in theory) or via cache-friendly scans of the original text (in practice).
@@
Paper Organization. In Section 2, we provide the necessary notation and tools. In Section 3, we provide an overview of previous related work. In Section 4, we present our framework, and in Section 5, we present our experimental evaluation. We conclude this paper in Section 6.
@@ 2 Preliminaries
In this section we provide useful background information to support subsequent descriptions.
@@ 2.1 Notation and Computational Model
Basic Notation. Recall that we deal with a string $T[0\mathinner{.\,.}n)$ of length $n$ over an alphabet $\Sigma=[0,\sigma)$. We assume that $\Sigma$ is an integer alphabet of polynomial size in $n$, i.e., $\sigma=n^{\mathcal{O}(1)}$. A substring of $T$ of length $k>0$ is called a $k$-mer. Given a query pattern $P[0\mathinner{.\,.}m)$ with $m\geq\ell$ for some lower bound $\ell>0$, the goal is to support $\textup{{Locate}}(P,T)=\{0\leq i<n-m+1\,|\,T[i\mathinner{.\,.}i+m)=P\}$. We refer to the number of occurrences of $P$ in $T$, that is, $|\textup{{Locate}}(P,T)|$, with $\textup{{Count}}(P,T)$.
The Computational Model. We assume that we have random read-only access to $T$ and count the space (in number of words) occupied on top of the space occupied by $T$. We assume the standard word RAM model of computation with machine words of $\Omega(\log n)$ bits.
@@ 2.2 Algorithmic Toolbox
The solutions we describe in this paper rely on few, well-defined, tools that we present below.
Minimizers. In this paper, we use a specific class of randomized methods to sketch a string, called minimizers [31, 32]. Minimizers are defined as the triple $(k,w,h)$: from a window of $w$ consecutive $k$-mers of $T$, the leftmost smallest $k$-mer according to the order $h$ is chosen as the minimizer of the window. Since at least one $k$-mer must be chosen every $w$ positions, the fraction of sampled $k$-mers — defined as the density of the sampling algorithm — is always at least $1/w$.
Several minimizer sampling algorithms have been proposed in the literature. See Section 3 of [18] for a recent overview of different sampling strategies and orders that lead to different densities. In this paper, however, we use the folklore random minimizer sampling, which is as defined above when $h$ is a pseudo-random hash function. We have the following result.
@@
In the context of this paper, we fix $\ell$ to be the minimum pattern length and let $w=\ell-k+1$. Each substring of length $\ell$ of $T$ therefore contains one minimizer. (In practice, we expect to have $|P|\gg\ell$ and that the sketch of $P$ is a sequence of several minimizers.) Further, we let $\mathcal{M}_{\ell,k}(T)$ indicate the sorted list of positions in $T$ of the minimizers of $T$. Let $z=|\mathcal{M}_{\ell,k}(T)|$ be the number of minimizers. By Theorem 2.1, we have that $z\approx 2/(\ell-k+2)$ in expectation (neglecting lower-order terms).
Tries. Given a set $\mathcal{X}$ of strings over the alphabet $\Sigma$, a trie $\textup{{Trie}}(\mathcal{X})$ is a rooted tree whose nodes represent the prefixes of the strings in $\mathcal{X}$. The edges of $\textup{{Trie}}(\mathcal{X})$ are labeled by letters from $\Sigma$; the prefix corresponding to node $u$ is denoted by $\textsf{str}(u)$ and is given by the concatenation of the letters labeling the path (sequence of edges) from the root of $\textup{{Trie}}(\mathcal{X})$ to $u$. The node $u$ is called the locus of $\textsf{str}(u)$. The parent-child relationship in $\textup{{Trie}}(\mathcal{X})$ is defined as follows: the root node is the locus of the empty string $\varepsilon$; and the parent $u$ of another node $v$ is the locus of $\textsf{str}(v)$ without the last letter. This letter is the edge label of $(u,v)$. The order on $\Sigma$ induces an order on the edges outgoing from any node of the trie. A node $u$ is branching if it has at least two children and terminal if $\textsf{str}(u)\in\mathcal{X}$.
A compacted trie is obtained from the underlying trie by removing all nodes except the root, the branching nodes, and the terminal nodes. More specifically, a compacted trie is a trie where all unary paths are collapsed into a single edge, labeled by the string obtained by concatenating all the letters of the edges in the unary path. The compacted trie takes $\mathcal{O}(|\mathcal{X}|)$ space provided that the edge labels are implicitly represented as pointers to fragments of strings in $\mathcal{X}$. Given the lexicographic order on $\mathcal{X}$ along with the lengths of the longest common prefixes between any two consecutive elements (in this order) of $\mathcal{X}$, one can compute $\textup{{Trie}}(\mathcal{X})$ in $\mathcal{O}(|\mathcal{X}|)$ time [25].
Rolling Hashing. Let $p$ be a prime number and choose $r\in[0,p)$ uniformly at random. The rolling hash value of $T[i\mathinner{.\,.}j]$ — which we simply refer to as fingerprint in the following — is defined as [24]:
@@
The adjective “rolling” refers to the way the hash value is updated incrementally as a fixed length window slides through the string $T$. The function $\phi$ allows to compute the fingerprint of a window just knowing the fingerprint of the previous window and the character that is being removed/added, instead of recalculating the fingerprint from scratch.
Clearly, if $T[i\mathinner{.\,.}i+\ell]=T[j\mathinner{.\,.}j+\ell]$ then $\phi(T[i\mathinner{.\,.}i+\ell])=\phi(T[j\mathinner{.\,.}j+\ell])$. On the other hand, if $T[i\mathinner{.\,.}i+\ell]\neq T[j\mathinner{.\,.}j+\ell]$ then $\phi(T[i\mathinner{.\,.}i+\ell])\neq\phi(T[j\mathinner{.\,.}j+\ell])$ with probability at least $1-\ell/p$ [9].
Since we are comparing only substrings of equal length, the number of different possible substring comparisons is less than $n^{3}$. Thus, for any constant $c\geq 1$, we can set $p$ to be a prime larger than $\max(|\Sigma|,n^{c+3})$ to make the function $\phi$ perfect (i.e., no collisions) with probability at least $1-n^{-c}$ (this means with high probability). Any fingerprint of $T$ or $P$ fits in one machine word, so that comparing any two fingerprints takes $\mathcal{O}(1)$ time. In particular, we will use the following well-known fact.
@@ 3 Related Work
Text indexing for matching long patterns (i.e., with lengths at least $\ell$ for some $l>0$) in the uncompressed setting has attracted some attention in the literature [8, 16, 26, 27, 2, 3]. The common idea of these approaches is to use some form of sketching, such as alphabet sampling [8], minimizer-like anchors [16, 26, 27, 2] or their worst-case counterparts [3]. The work of [8] chooses a subset of the alphabet and constructs a sparse suffix array only for the suffixes starting with a letter from the chosen subalphabet. The search starts with finding the leftmost occurrence $j$ of a sampled letter of pattern $P$. Then the suffix $P[j\mathinner{.\,.}m)$ is sought using the sparse suffix array with standard means. After that, each occurrence of the suffix is verified against the text with the previous $j-1$ letters. The work of [16] proposes a similar approach. It first computes the set $B$ of starting positions of the minimizers of text $T$ and then constructs the sparse suffix array only for the suffixes starting at the positions in $B$. Upon a query pattern $P$, it computes the starting position $j$ of the leftmost minimizer of $P$, thus implicitly partitioning $P$ into $P[0\mathinner{.\,.}j)$ and $P[j\mathinner{.\,.}m)$. It then searches $P[j\mathinner{.\,.}m)$ in the sparse suffix array, and verifies each occurrence of it using letter comparisons against $T$ using $P[0\mathinner{.\,.}j)$. Subsequent works [26, 27, 2] propose to also construct a sparse suffix array for the reversed prefixes ending at the positions in $B$, and conceptually link the two suffix arrays with a geometric data structure. As opposed to [8, 16], these approaches [26, 27, 2] thus offer query times with theoretical guarantees. An important practical limitation of these works is that they rely on sparse suffix sorting which is a rather undeveloped topic in practical terms [4]. From a theory perspective, the following is known.
@@
For any string $T$ of length $n$ over an integer alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$ and any integer $\ell>0$, we can construct an index that occupies $\mathcal{O}(n/\ell)$ extra space and reports all $\textup{{Count}}(P,T)$ occurrences of any pattern $P$ of length $|P|\geq\ell$ in $T$ in $\mathcal{\tilde{O}}(|P|+\textup{{Count}}(P,T))$ time. The index can be constructed in $\mathcal{\tilde{O}}(n)$ time and $\mathcal{O}(n/\ell)$ working space.
The practical limitation of Theorem 3.1 is that it relies on an intricate sampling scheme and on geometric data structures which are both unlikely to be efficient in practice; see [3] for more details.
Another common characteristic of the aforementioned approaches is that they are not universal. They enhance the text with specific data structures (typically, the sparse suffix array of the sampled suffixes and some geometric data structures) and so they also have a specific query algorithm. The main benefit of the approach we describe in this paper is that it can be used with (and improve) any text indexing technique.
Other Related Work. There also exists work [21] that attempts to accelerate indexing lookup by working in sketch space (in this case, using a prefix-free parse [5] of the text and pattern). This approach builds an index over both the original and sketched text, however, and has been explored only in the context of compressed indexes (i.e., the FM-index).
Figure 1: The U-index framework. The steps (1) and (2) are to build the index. The steps (3)–(6) are to query with the framework. The sketching scheme in steps (1) and (3) must be the same.

### v3

@@ Abstract
Motivation.
Text indexing is a fundamental and well-studied problem. Classic solutions to this problem either replace the original text with a compressed representation, e.g., the FM-index and its variants, or keep it uncompressed but attach some redundancy — an index — to accelerate matching, e.g., the suffix array. The former solutions thus retain excellent compressed space, but are practically slow to construct and query. The latter approaches, instead, sacrifice space efficiency but are typically faster; for example, the suffix array takes much more space than the text itself for commonly used alphabets, like ASCII or DNA, but it is very fast to construct and query.
Methods.
In this paper, we show that efficient text indexing can be achieved using just a small extra space on top of the original text, provided that the query patterns are sufficiently long. More specifically, we develop a new indexing paradigm in which a sketch of a query pattern is first matched against a sketch of the text. Once candidate matches are retrieved, they are verified using the original text. This paradigm is thus universal in the sense that it allows us to use any solution to index the sketched text, like a suffix array, FM-index, or r-index.
Results.
We explore both the theory and the practice of this universal framework. With an extensive experimental analysis, we show that, surprisingly, universal indexes can be constructed much faster than their unsketched counterparts and take a fraction of the space, as a direct consequence of (i) having a lower bound on the length of patterns and (ii) working in sketch space. Furthermore, these data structures have the potential of retaining or even improving query time, because matching against the sketched text is faster and verifying candidates can be theoretically done in constant time per occurrence (or, in practice, by short and cache-friendly scans of the text).
Finally, we discuss some important applications of this novel indexing paradigm to computational biology. We hypothesize that such indexes will be particularly effective when the queries are sufficiently long, and so we demonstrate applications in long-read mapping.
@@ Results.
Text Indexing; Sketching; Minimizers; Hashing
@@ 1 Introduction
Given a string $T[0\mathinner{.\,.}n)$ (the “text” henceforth) of $n$ characters over an alphabet $\Sigma=[0,\sigma)$, we are asked to pre-process $T$ so that the following queries are supported efficiently for any string $P[0\mathinner{.\,.}m)$ (the “pattern” henceforth) and any $0\leq i<j\leq n$:
@@
On the other hand, uncompressed approaches, such as the suffix array [31], do not alter the original text but instead attach some form of redundancy — henceforth referred to as the index — to accelerate pattern matching. These methods, while being faster in terms of query times, often suffer from significant space inefficiency. For example, suffix arrays usually take up much more space than the text itself, particularly for commonly used alphabets such as ASCII or DNA [21]. This makes such structures impractical for many applications, especially when working with very large datasets.
Our Contributions.
In this paper, we focus on the latter approach of uncompressed text indexing and aim to address its space inefficiency while retaining efficient queries. In the following, we discuss the Locate query only, given that Count can be trivially implemented using Locate, and Extract is done by explicitly accessing the text $T$. We propose a novel approach that solves the text indexing problem using only a small amount of additional space beyond the original text, without sacrificing the speed of query processing. This is achieved under the assumption that the query patterns $P[0\mathinner{.\,.}m)$ are sufficiently long, i.e., we require that $m\geq\ell$ for some fixed lower bound $\ell>0$ [3], and operate under the assumption that $n\gg\ell$. A typical value for $\ell$ lies in $[32,1000]$. The core idea of our approach is to transform the text into sketch space, and to construct and use a smaller index over the sketched text to enable queries with only small additional space. Generally speaking, a sketch is a compact representation of an object that retains enough information for approximate matches.
@@
    Each query pattern $P$ is then sketched into $Q=\textup{{Sketch}}(P)$, and $Q$ is matched against $\textup{{Index}}(S)$ to identify potential matches $L^{\prime}=\textup{{Locate}}(Q,S)$.
@@
    Candidate matches in $L^{\prime}$ are mapped back to their positions in $T$, where we verify that indeed $P$ matches $T$. Thus we obtain $L=\textup{{Map}}(L^{\prime})$, where $L=\textup{{Locate}}(P,T)$.
The universality of our framework lies in its flexibility: (i) any indexing structure, Index, such as the FM-index or the suffix array, can be used to index the sketched text $S$, making the method adaptable to a variety of text indexing techniques [31, 14, 20, 15, 27, 28]; and (ii) any locally consistent sampling mechanism, Sketch, such as minimizers [34, 35], syncmers [11], or bd-anchors [30], can be used to sketch $S$ and $P$, making the method adaptable to a variety of sampling techniques. We also expect that our framework, which relies on $\textup{{Index}}(S)$, has competitive queries compared to the ones of $\textup{{Index}}(T)$ or possibly faster when the query time of $\textup{{Index}}(T)$ is a function of $n=|T|$. This is merely because $|S|<|T|$ (in practice, depending on the sketching mechanism used, $|S|$ could be $\Theta(|T|/\ell)$ or $\Theta(|T|/(\ell-k))$ for some $k=o(\ell)$). A typical such case is when Index is the suffix array [31] and binary search will take place on a smaller array, i.e., the suffix array of $S$ instead of the suffix array of $T$.
We explore both the theoretical underpinnings and the practical implementation of this universal framework. We demonstrate that universal indexes can be constructed significantly faster and occupy a fraction of the space compared to their unsketched counterparts. These space savings are a direct consequence of operating in the reduced “sketch space”. Depending on the index used, the performance of pattern matching in sketched space can be either faster (suffix array) or slower (FM-index). In either case, the overall query performance is slightly slower though, due to a relatively large number of false positive matches that have to be rejected during verification. This verification step can be performed in constant time per occurrence (in theory) or via cache-friendly scans of the original text (in practice).
@@
Paper Organization.
In Section 2, we provide the necessary notation and tools. In Section 3, we provide an overview of previous related work. In Section 4, we present our framework, and in Section 5, we present our experimental evaluation. We conclude this paper in Section 6.
@@ 2 Preliminaries
In this section, we provide useful background information to support subsequent descriptions.
@@ 2.1 Notation and Computational Model
Basic Notation.
Recall that we deal with a string $T[0\mathinner{.\,.}n)$ of length $n$ over an alphabet $\Sigma=[0,\sigma)$. We assume that $\Sigma$ is an integer alphabet of polynomial size in $n$, i.e., $\sigma=n^{\mathcal{O}(1)}$. A substring of $T$ of length $k>0$ is called a $k$-mer. Given a query pattern $P[0\mathinner{.\,.}m)$ with $m\geq\ell$ for some lower bound $\ell>0$, the goal is to support the following operation:
$\textup{{Locate}}(P,T)=\{0\leq i<n-m+1\,|\,T[i\mathinner{.\,.}i+m)=P\}.$
We refer to the number of occurrences of $P$ in $T$, that is, $|\textup{{Locate}}(P,T)|$, with $\textup{{Count}}(P,T)$.
The Computational Model.
We assume that we have random read-only access to $T$ and count the space (in number of words) occupied on top of the space occupied by $T$. We assume the standard word RAM model of computation with machine words of $\Omega(\log n)$ bits.
@@ 2.2 Algorithmic Toolbox
The solutions we describe in this paper rely on few, well-defined tools that we present below.
Minimizers.
We use a specific class of randomized methods to sketch a string, called minimizers [34, 35]. Minimizers are defined as the triple $(k,w,h)$: from a window of $w$ consecutive $k$-mers of $T$, the leftmost smallest $k$-mer according to the order $h$ is chosen as the minimizer of the window. Since at least one $k$-mer must be chosen every $w$ positions, the fraction of sampled $k$-mers — defined as the density of the sampling algorithm — is always at least $1/w$. Several minimizer sampling algorithms have been proposed in the literature. See Section 3 of [19] for a recent overview of different sampling strategies and orders that lead to different densities. Here, however, we use the folklore random minimizer sampling, which is as defined above and uses a pseudo-random hash function for the order $h$. We have the following result.
@@
We fix $\ell$ to be the minimum pattern length and let $w=\ell-k+1$. Each substring of length $\ell$ of $T$ therefore contains one minimizer. (In practice, we expect to have $|P|\gg\ell$ and that the sketch of $P$ is a sequence of several minimizers.) Further, we let $\mathcal{M}_{\ell,k}(T)$ indicate the sorted list of positions in $T$ of the minimizers of $T$. Let $z=|\mathcal{M}_{\ell,k}(T)|$ be the number of minimizers. By Theorem 2.1, we have that $z\approx 2(|T|-\ell+1)/(\ell-k+2)$ in expectation.
Tries.
Given a set $\mathcal{X}$ of strings over the alphabet $\Sigma$, a trie $\textup{{Trie}}(\mathcal{X})$ is a rooted tree whose nodes represent the prefixes of the strings in $\mathcal{X}$. The edges of $\textup{{Trie}}(\mathcal{X})$ are labeled by letters from $\Sigma$; the prefix corresponding to node $u$ is denoted by $\textsf{str}(u)$ and is given by the concatenation of the letters labeling the path (sequence of edges) from the root of $\textup{{Trie}}(\mathcal{X})$ to $u$. The node $u$ is called the locus of $\textsf{str}(u)$. The parent-child relationship in $\textup{{Trie}}(\mathcal{X})$ is defined as follows: the root node is the locus of the empty string $\varepsilon$; and the parent $u$ of another node $v$ is the locus of $\textsf{str}(v)$ without the last letter. This letter is the edge label of $(u,v)$. The order on $\Sigma$ induces an order on the edges outgoing from any node of the trie. A node $u$ is branching if it has at least two children and terminal if $\textsf{str}(u)\in\mathcal{X}$.
A compacted trie is obtained from $\textup{{Trie}}(\mathcal{X})$ by removing all nodes except the root, the branching nodes, and the terminal nodes. A compacted trie is thus a trie where all unary paths are collapsed into a single edge, labeled by the string obtained by concatenating all the letters of the edges in the unary path. The compacted trie takes $\mathcal{O}(|\mathcal{X}|)$ space provided that the edge labels are implicitly represented as pointers to fragments of strings in $\mathcal{X}$. Given the lexicographic order on $\mathcal{X}$ along with the lengths of the longest common prefixes between any two consecutive elements (in this order) of $\mathcal{X}$, one can compute $\textup{{Trie}}(\mathcal{X})$ in $\mathcal{O}(|\mathcal{X}|)$ time [26].
Rolling Hashing.
Let $p$ be a prime number and choose $r\in[0,p)$ uniformly at random. The rolling hash value of $T[i\mathinner{.\,.}j]$ — which we term fingerprint — is defined as [25]:
@@
The adjective “rolling” refers to the way the hash value is updated incrementally as a fixed-length window slides through the string $T$. The function $\phi$ allows one to compute the fingerprint of a window just knowing the fingerprint of the previous window and the character that is being removed/added, instead of recalculating the fingerprint from scratch.
By definition, if $T[i\mathinner{.\,.}i+\ell]=T[j\mathinner{.\,.}j+\ell]$, then $\phi(T[i\mathinner{.\,.}i+\ell])=\phi(T[j\mathinner{.\,.}j+\ell])$. On the other hand, if $T[i\mathinner{.\,.}i+\ell]\neq T[j\mathinner{.\,.}j+\ell]$, then $\phi(T[i\mathinner{.\,.}i+\ell])\neq\phi(T[j\mathinner{.\,.}j+\ell])$ with probability at least $1-\ell/p$ [10]. Since we are comparing only substrings of equal length, the number of different possible substring comparisons is less than $n^{3}$. Thus, for any constant $c\geq 1$, we can set $p$ to be a prime larger than $\max(|\Sigma|,n^{c+3})$ to make the function $\phi$ perfect (i.e., no collisions) with probability at least $1-n^{-c}$ (this means with high probability). Any fingerprint of $T$ or $P$ fits in one machine word, so that comparing any two fingerprints takes $\mathcal{O}(1)$ time. In particular, we will use the following well-known fact.
@@ 3 Related Work
Text indexing for matching long patterns (i.e., with lengths at least $\ell$ for some $\ell>0$) in the uncompressed setting has attracted some attention in the literature [9, 17, 29, 30, 3, 4]. The common idea of these approaches is to use some form of sketching, such as alphabet sampling [9], minimizer-like anchors [17, 29, 30, 3] or their worst-case counterparts [4]. The work of [9] chooses a subset of the alphabet and constructs a sparse suffix array only for the suffixes starting with a letter from the chosen subalphabet. The search starts with finding the leftmost occurrence $j$ of a sampled letter of pattern $P$. Then the suffix $P[j\mathinner{.\,.}m)$ is sought using the sparse suffix array with standard means. After that, each occurrence of the suffix is verified against the text with the previous $j$ letters. The work of [17] proposes a similar approach. It first computes the set $B$ of starting positions of the minimizers of text $T$ and then constructs the sparse suffix array only for the suffixes starting at the positions in $B$. Upon a query pattern $P$, it computes the starting position $j$ of the leftmost minimizer of $P$, thus implicitly partitioning $P$ into $P[0\mathinner{.\,.}j)$ and $P[j\mathinner{.\,.}m)$. It then searches $P[j\mathinner{.\,.}m)$ in the sparse suffix array, and verifies each occurrence of it using letter comparisons against $T$ using the previous $j$ letters. Subsequent works [29, 30, 3] propose to also construct a sparse suffix array for the reversed prefixes ending at the positions in $B$, and conceptually link the two suffix arrays with a geometric data structure. As opposed to [9, 17], these approaches [29, 30, 3] thus offer query times with theoretical guarantees. An important practical limitation of these works is that they rely on sparse suffix sorting which is a rather undeveloped topic in practical terms [5]. From a theory perspective, the following is known.
@@
For any string $T=T[0\mathinner{.\,.}n)$ over an alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$ and any integer $\ell>0$, we can construct an index that occupies $\mathcal{O}(n/\ell)$ extra space and reports all $\textup{{Count}}(P,T)$ occurrences of any pattern $P$ of length $|P|\geq\ell$ in $T$ in $\mathcal{\tilde{O}}(|P|+\textup{{Count}}(P,T))$ time. The index can be constructed in $\mathcal{\tilde{O}}(n)$ time and $\mathcal{O}(n/\ell)$ working space.
The practical limitation of Theorem 3.1 is that it relies on an intricate sampling scheme and on geometric data structures which are both unlikely to be efficient in practice [4].
Another common characteristic of the aforementioned approaches is that they are not universal. They enhance the text with specific data structures (typically, the sparse suffix array of the sampled suffixes and some geometric data structures) and so they also have a specific query algorithm. The main benefit of the approach we describe in this paper is that it can be used with (and improve) any text indexing technique [31, 14, 20, 15, 27, 28].
Other Related Work.
There also exists work [22] that attempts to accelerate indexing lookup by working in sketch space (in this case, using a prefix-free parse [6] of the text and pattern). This approach, however, builds an index over both the original and the sketched text, and has been explored only in the context of compressed indexes (i.e., the FM-index).